# LLM-Based Agentic Systems for Software Engineering: Challenges and Opportunities

### v1

@@ 1 Introduction
Large Language Models (LLMs) [62] have shown promising potential to transform the Software Development Lifecycle (SDLC) . Their exceptional capabilities in text understanding, knowledge inference, and reasoning have opened new avenues for addressing complex software engineering tasks. Both academia and industry have explored a wide range of techniques to address various Software Engineering (SE) use cases, such as prompt engineering [38, 44], in-context learning [9, 45], parameter-efficient fine-tuning [15], and retrieval augmented generation [13]. Each approach offers distinct advantages for tailoring LLM capabilities to specific SE tasks.
@@ 3.2 Code Generation
Code generation with iterative refinement: Multi-agent systems can dynamically refine the generated code through iterative feedback from models [51, 26, 4, 20], tools [60, 54, 61, 60, 16, 51, 41, 48], and humans [49, 34]. For instance, the agent can receive peer-reflection from multiple models with specialized roles [51, 26] or conduct self-refinement based on the previous output [4, 20]. The agent can also collect informative compilation errors from compiler and interpreter tools [60, 54], or even get access to external resources by applying retrieval or searching tools [61, 60, 16]. Furthermore, hybrid feedback from tool and model is possible – receiving error messages from tools and then using LLMs’ contextual understanding to provide corresponding explanations, suggestions, and instructions [51, 41, 48]. Considering the critical role of human in clarifying ambiguous requirements, several works incorporate human feedback into the code generation process [49, 34].

### v2

@@ 1 Introduction
Large Language Models (LLMs) [64] have shown promising potential to transform the Software Development Lifecycle (SDLC) . Their exceptional capabilities in text understanding, knowledge inference, and reasoning have opened new avenues for addressing complex software engineering tasks. Both academia and industry have explored a wide range of techniques to address various Software Engineering (SE) use cases, such as prompt engineering [39, 45], in-context learning [9, 46, 47], parameter-efficient fine-tuning [16], and retrieval augmented generation [14]. Each approach offers distinct advantages for tailoring LLM capabilities to specific SE tasks.
@@ 3.2 Code Generation
Code generation with iterative refinement: Multi-agent systems can dynamically refine the generated code through iterative feedback from models [53, 27, 4, 21], tools [62, 56, 63, 62, 17, 53, 42, 50, 10], and humans [51, 35]. For instance, the agent can receive peer-reflection from multiple models with specialized roles [53, 27] or conduct self-refinement based on the previous output [4, 21]. The agent can also collect informative compilation errors from compiler and interpreter tools [62, 56], or even get access to external resources by applying retrieval or searching tools [63, 62, 17]. Furthermore, hybrid feedback from tool and model is possible – receiving error messages from tools and then using LLMs’ contextual understanding to provide corresponding explanations, suggestions, and instructions [53, 42, 50]. Considering the critical role of human in clarifying ambiguous requirements, several works incorporate human feedback into the code generation process [51, 35].
@@ Acknowledgments
  • [10] H. Elhashemy, Y. Lotfy, and Y. Tang (2025) Bridging the prototype-production gap: a multi-agent system for notebooks transformation. arXiv preprint arXiv:2511.07257. Cited by: §3.2.
@@
  • [47] N. N. Vaidya, T. A. Runkler, T. Hubauer, V. Haderlein-Hoegberg, and M. M. Brandt (2025) Conceptual in-context learning and chain of concepts: solving complex conceptual problems using large language models. In 2025 IEEE Symposium on Computational Intelligence in Natural Language Processing and Social Media (CI-NLPSoMe), Vol. , pp. 1–7. External Links: Document Cited by: §1.